OpenStudy (surryyy):

If we toss a pair of dice, what is the probability that a sum of 7 will appear? Answer 1/36 7/36 1/2 1/6

OpenStudy (gorv):

how 7 come on two dice??

OpenStudy (gorv):

when sum of 7 will appear?? condition??

OpenStudy (surryyy):

i dont understand.

OpenStudy (gorv):

hmm...like one number on first dice and other on second when their sum became 7

OpenStudy (gorv):

hmm..like 1+6=7 like this what are the combination that produce sum as seven

OpenStudy (surryyy):

oh right! 2+5 3+4 4+3 5+2 6+1

OpenStudy (akashdeepdeb):

Find the Sample Space!

OpenStudy (surryyy):

5/36?

ganeshie8 (ganeshie8):

thats a good try, but looks like you have missed one favorable outcome : 1+6

OpenStudy (akashdeepdeb):

2,5 + 5,2 + 6,1 + 1,6 + 3,4 + 4,3. [6/36 = 1/6] :)

OpenStudy (surryyy):

Thnx! i understand!